# Reinforcing Clinical Decision Support through Multi-Agent Systems and Ethical AI Governance

### v1

@@ I Introduction
AI has steadily made its way into many industries, from image recognition and supply chain logistics [1, 2, 3], writing assist [4] to the field of healthcare [5, 6]. Within medicine, it’s proving valuable for sharpening diagnostic precision, supporting treatment planning, and helping clinicians keep a closer eye on patients [7]. A growing body of work has focused on using AI to interpret complex medical visuals like surgical footage [8], X-rays [9], CTs [10], and MRI scans [11], making interpretation faster and more consistent. But it doesn’t stop with images—AI is also being used to make sense of physiological signals, including EEG [12, 13], ECG [14], and data from wearable sensors [15, 16]. These efforts are opening up new possibilities across neurology, psychiatry, and continuous patient monitoring [17]. Altogether, these advancements point to a future where AI supports both visual and signal-based insights, forming the backbone of smarter clinical decision-making tools.
Clinical decision support systems (CDSS) have become a vital part of today’s healthcare settings, offering insights drawn from electronic health records (EHRs) and real-time monitoring tools. Yet, many of the traditional AI models used in these systems fall short when it comes to flexibility, transparency, and oversight—key qualities especially critical in high-risk settings like intensive care units (ICUs). To address these limitations, we introduce a modular multi-agent system (MAS) designed to reflect how clinical teams make decisions, with a built-in emphasis on ethical AI governance to uphold both safety and accountability.
Building on progress in agent-based frameworks and the coordinated use of large language models (LLMs), our system breaks down the decision-making pipeline into focused, collaborative agents. Each one is responsible for a different aspect of ICU assessment: from interpreting lab results and tracking vital signs to making context-sensitive judgments based on a patient’s history or co-existing conditions. These individual agents pass their findings to a central integration agent that brings everything together, enabling more comprehensive predictions and cross-validated outcomes. This structure mimics how doctors work—gathering evidence from various sources, weighing context, and forming a unified clinical picture.
@@ II-A Applications of Clinical Decision Support Systems in Intensive Care Settings
Clinical Decision Support Systems (CDSS) have come a long way, especially in ICU environments where every second counts. Earlier systems typically leaned on rule-based logic or statistical methods to generate recommendations [19, 20]. More recent developments have looked to Clinical Practice Guidelines (CPGs) as a way to enrich Large Language Models (LLMs), boosting their ability to offer context-aware treatment advice. Research suggests that LLMs enhanced with CPGs outperform traditional models in delivering more accurate clinical suggestions [19].
@@
The eICU Collaborative Research Database has become a cornerstone for ICU-related studies, offering detailed information from more than 200,000 ICU stays across the United States [21]. With data ranging from vital signs and treatment plans to severity scores, diagnoses, and interventions, it provides a rich foundation for building and evaluating AI models tailored to critical care environments.
The electronic ICU (eICU) itself is a telemedicine-driven model aimed at delivering continuous critical care for high-risk patients [22]. Systems like Philips’ eICU extend ICU expertise to the bedside through the eCareManager platform, which connects hospital networks and generates real-time clinical insights [23].
Deploying eICU systems has led to tangible benefits. At Baptist Health South Florida, for example, implementation was followed by a 23% drop in ICU mortality rates and up to a 25% reduction in patient length of stay [23].
@@ II-C LLM-Based Agents in Healthcare
The rapid advancement of Large Language Models (LLMs) has led to their growing presence in the healthcare space. These models now power a variety of applications—ranging from virtual assistants and personalized health education to symptom checkers and mental health support tools [24]. By streamlining both patient care and administrative tasks, LLM-based agents are poised to reshape how services are delivered.
One notable example is MDAgents, a multi-agent system built on LLMs that tackles complex medical decision-making tasks [25]. Its architecture is designed to reflect how healthcare teams operate in real-world settings, enabling dynamic collaboration among agents. Performance benchmarks show that MDAgents outperforms earlier models across multiple evaluation criteria.
A recent review takes a deep dive into the evolving landscape of LLM-driven agents in medicine [26], covering their technical underpinnings, practical uses, and ongoing challenges. The survey highlights key system elements like planning mechanisms, reasoning strategies, tool integrations, and design configurations. These agents are already being applied in areas such as clinical support, documentation automation, simulation training, and workflow optimization.
@@ II-D Multi-Agent Systems in Healthcare
II-E Ethical Governance and Explainability in Healthcare AI
@@ II-F Motivation and Research Gap
While clinical decision support systems, LLM-powered agents [36], and multi-agent frameworks have made notable strides in healthcare, there’s still a considerable gap when it comes to integrating these technologies into real-world ICU settings under clear ethical oversight. Many current solutions fall short in modularity, lack transparency, or aren’t built with the kind of structured, inter-agent communication needed to reflect the fast-paced, interdisciplinary nature of intensive care.
@@ III-A Dataset and Preprocessing
We based our study on the eICU Collaborative Research Database v2.0 [21], which provides a wide range of de-identified ICU data from over 200,000 admissions across various U.S. hospitals. This dataset combines both structured information—like vital signs and lab results—and unstructured clinical notes, including nursing and physician documentation, offering a comprehensive view of each patient’s stay.
To build our pipeline, we worked directly with a set of key tables from the eICU database, including patient.csv.gz, lab.csv.gz, vitalPeriodic.csv.gz, note.csv.gz, and medication.csv.gz, along with APACHE-related files like apacheApsVar.csv.gz and apachePatientResult.csv.gz. We grouped records by patientunitstayid to keep everything aligned per ICU stay. Incomplete entries—such as patients without enough data in labs, vitals, or notes—were excluded to ensure each case provided a meaningful input for modeling.
The preprocessing wasn’t overly complicated but required care. We cleaned up missing values, sorted events based on when they happened (using time offsets), and trimmed overly long text fields so they’d fit within the token limits of our language model components. The goal here wasn’t just to make the data model-ready, but to keep as much useful clinical context as possible without overwhelming the system.
From the full dataset, we extracted several types of information per patient to construct meaningful input features. For vital signs, we collected the latest ten time-ordered readings per patient to preserve recent physiological trends. In the laboratory domain, we selected the most recent non-redundant values for a set of clinically relevant biomarkers. To capture context from unstructured text, we sampled up to three representative clinical notes per patient, focusing on physician and nursing documentation.
For treatments and medications, we focused on the top 20 items either by uniqueness or frequency across the cohort. Lastly, APACHE scores and predicted outcomes were included as reference metrics for downstream validation and benchmarking.
@@ III-B Multi-Agent Framework Design
  • Vitals Analysis Agent: Processes vital signs (e.g., HR, SBP, SpO2, temperature) and evaluates physiological stability, respiratory function, and cardiovascular performance.
@@
Each agent is implemented using Anthropic Claude 3.7 Sonnet and configured via the intelli.flow framework, which allows asynchronous agent orchestration with JSON-structured prompts and logging.
@@ III-D System Execution and DAG Orchestration
Execution is managed asynchronously using Python’s asyncio to allow concurrent LLM calls and reduce latency. Flow state is logged per patient and visualized via networkx-based flow diagrams saved as PNGs. The system supports multi-threaded batch evaluation and error-tolerant retries.
III-E Implementation and Agent API Configuration
Each agent is instantiated as a Claude-based text agent with a predefined mission, API credentials, and output format enforcement. Prompts are customized per task using structured sections (e.g., “KEY ABNORMALITIES”, “APACHE RELEVANT FINDINGS”). Inputs are truncated or summarized to fit within the 10,000-token limit of Claude.
@@ III-E Implementation and Agent API Configuration
  model_params={"key": ANTHROPIC_API_KEY,
  "model": "claude-3-7-sonnet-20250219"}
@@ III-F Ethical AI Governance and Explainability
III-G Quantitative the Agent Transparency
The TransparencyMetrics [37] class evaluates the transparency of clinical prediction explanations by analyzing text responses for key transparency features. It calculates a transparency score by checking for the presence of five critical elements: explicit weights, monotonic relationships, feature importance, decision path explanations, and uncertainty quantification. The scoring process uses regular expression pattern matching to detect relevant terms in the response content. The final transparency score is normalized by dividing the raw score (number of detected features) by the maximum possible score (total number of transparency features), resulting in a value between 0 and 1 that quantifies how transparent and interpretable a clinical prediction explanation is.
@@ IV Results
Figure 2: Comparison of performance between the Multi-agent and Single-agent frameworks across three evaluation metrics: Mortality Accuracy, Length of Stay (LOS) Mean Error, and LOS Median Error. Each model was executed 10 times, and the box plots represent the distribution of results over these runs. The Multi-agent framework shows slightly higher mean mortality accuracy with slightly more variance, while LOS-related errors are nearly identical between the two models. These results indicate comparable predictive performance, with minor differences in consistency and central tendency.
The experimental results in Table I and Figure 2 indicate that both the multi-agent and single-agent approaches yield remarkably similar performance across all evaluated metrics. Each experiment was conducted ten times per setting, and the reported values represent the average across these ten runs to ensure consistency and robustness of evaluation. In terms of mortality prediction accuracy, the multi-agent model achieved a mean of 85.54%, while the single-agent model reached 85.52%. Although the multi-agent system shows a slightly higher mean, the difference of 0.02% is minimal and likely not statistically significant. Additionally, the standard deviation for the multi-agent model is marginally higher, suggesting a bit more variability across runs.
TABLE I: Comparison of Multi-agent and Single-agent Models (Average over 10 runs)
Metric Model Mean Std
Mortality Accuracy (%) Multi-agent 85.543 0.00088
Single-agent 85.520 0.00075
LOS Mean Error (days) Multi-agent 0.74911 0.00107
Single-agent 0.74906 0.00092
LOS Median Error (days) Multi-agent 0.74987 0.00167
Single-agent 0.74882 0.00147
When evaluating the prediction of length of stay (LOS), both the mean and median error metrics show almost no practical difference between the two approaches. The mean LOS error is identical for both models at approximately 0.7491 days, and the median LOS error differs only slightly—0.7499 days for the multi-agent model versus 0.7488 days for the single-agent. Again, the single-agent model demonstrates slightly lower variance, but the magnitude of the difference remains negligible.
TABLE II: Comparison of transparency score of Multi-agent and Single-agent Models (Average over 10 runs)
@@
Average transparency score (%) Multi-agent 60.01
Single-agent 75.04
Table II presents a comparison of the average transparency scores achieved by the multi-agent and single-agent models across 10 independent runs. The results indicate a clear advantage for the single-agent approach, which attained a significantly higher average transparency score of 75.04%, compared to 60.01% for the multi-agent model. This suggests that the single-agent architecture produces more interpretable or explainable outputs under the current evaluation framework. One possible explanation is that, without explicit inter-agent communication or coordination strategies, the multi-agent model’s internal decision-making process may become more distributed and opaque, reducing its overall transparency. In contrast, the centralized structure of the single-agent model may allow for more consistent and traceable reasoning paths, leading to better transparency metrics.
@@
Overall, the performance of the multi-agent and single-agent models is effectively equivalent in terms of predictive accuracy and error, with only marginal differences that do not clearly favor one approach over the other. Interestingly, in this case, the single-agent model shows slightly better results. This may be due to the absence of explicit interaction mechanisms between agents, which limits the potential benefits of distributed learning. Without proper coordination or communication strategies, the multi-agent system may not fully leverage its structure, making centralized processing by a single agent more effective. This is an insightful finding that highlights how, in the absence of well-designed inter-agent dynamics, simpler single-agent architectures can outperform more complex multi-agent frameworks. Unless there are additional considerations—such as scalability, modularity, or deployment constraints—that specifically favor the multi-agent setup, either model could be considered a valid choice based on the observed outcomes.
@@ V Conclusion
In this study, we compared the performance of single-agent and multi-agent models within a predictive learning framework. Our results show that both approaches achieve comparable levels of predictive accuracy and error, with only marginal differences observed. Interestingly, the single-agent model consistently performed slightly better. We hypothesize that this may be due to the absence of explicit interaction or coordination mechanisms between agents in the current multi-agent setup. Without such mechanisms, the benefits of distributed learning cannot be fully realized, and centralizing all information within a single agent may lead to more effective learning.
This finding highlights an important consideration: multi-agent architectures do not inherently guarantee improved performance unless supported by well-designed inter-agent communication or collaboration strategies. Therefore, unless future system requirements demand advantages such as scalability, modularity, or robustness—areas where multi-agent systems typically excel—the simpler single-agent design may be a more practical and effective choice under current conditions.

### v2

@@ I Introduction
Artificial intelligence (AI) has steadily made its way into many industries, from image recognition and supply chain logistics [1, 2, 3], writing assist [4] to the field of healthcare [5, 6]. Within medicine, it’s proving valuable for sharpening diagnostic precision, supporting treatment planning, and helping clinicians keep a closer eye on patients [7]. A growing body of work has focused on using AI to interpret complex medical visuals like surgical footage [8], X-rays [9], computed tomography (CT) scans [10], and magnetic resonance imaging (MRI) scans [11], making interpretation faster and more consistent. But it doesn’t stop with images. AI is also being used to make sense of physiological signals, including electroencephalography (EEG) [12, 13], electrocardiography (ECG) [14], and data from wearable sensors [15, 16]. These efforts are opening up new possibilities across neurology, psychiatry, and continuous patient monitoring [17]. Altogether, these advancements point to a future where AI supports both visual and signal-based insights, forming the backbone of smarter clinical decision-making tools.
Clinical decision support systems (CDSS) have become a vital part of today’s healthcare settings, offering insights drawn from electronic health records (EHRs) and real-time monitoring tools. Yet, many of the traditional AI models used in these systems fall short when it comes to flexibility, transparency, and oversight key qualities, especially critical in high-risk settings like intensive care units (ICUs). To address these limitations, we introduce a modular multi-agent system (MAS) designed to reflect how clinical teams make decisions, with a built-in emphasis on ethical AI to uphold both explainability and accountability.
Building on progress in agent-based frameworks and the coordinated use of large language models (LLMs), our system breaks down the decision-making pipeline into focused, collaborative agents. Each one is responsible for a different aspect of ICU assessment: from interpreting lab results and tracking vital signs to making context-sensitive judgments based on a patient’s history or co-existing conditions. These individual agents pass their findings to a central integration agent that brings everything together, enabling more comprehensive predictions and cross-validated outcomes. This structure mimics how doctors gathering evidence from various sources, weighing context, and forming a unified clinical picture.
@@ II-A Applications of Clinical Decision Support Systems in Intensive Care Settings
Clinical decision support systems (CDSS) have come a long way, especially in ICU environments where every second counts. Earlier systems typically leaned on rule-based logic or statistical methods to generate recommendations [19, 20]. More recent developments have looked to clinical practice guidelines (CPGs) as a way to enrich LLMs, boosting their ability to offer context-aware treatment advice. Research suggests that LLMs enhanced with CPGs outperform traditional models in delivering more accurate clinical suggestions [19].
@@
The eICU Collaborative Research Database has emerged as a critical resource for intensive care research, gathering comprehensive data from over 200,000 ICU stays across the United States [21]. This extensive collection spans vital parameters—including vital signs, treatment protocols, severity indices, diagnoses, and interventions—serving as a solid groundwork for developing and validating AI models that address the specific challenges of critical care.
The electronic ICU relies on a telemedicine framework to monitor high-risk patients and deliver critical care, even when onsite specialists are not available [22]. For instance, Philips has developed an eICU system that utilizes the eCareManager platform to bring expert ICU support directly to the patient’s bedside. By linking hospital networks and supplying real-time clinical feedback, this approach narrows the gap between off-site experts and immediate patient needs [23].
In practical terms, the rollout of eICU systems has produced measurable improvements in patient outcomes. At Baptist Health South Florida, the introduction of the eICU model was linked to a 23% decrease in ICU mortality and a reduction in the average length of stay by up to 25% [23]. These results not only demonstrate how telemedicine can streamline critical care delivery but also highlight the potential for such systems to transform clinical practice through enhanced real-time decision-making and optimized resource use.
@@ II-C LLM-Based Agents in Healthcare
Large language models (LLMs) have recently become more common in healthcare. These models now assist in multiple areas, including virtual assistants, individualized health education, symptom checking, and mental health support tools [24]. By improving patient interactions and simplifying administrative tasks, LLM-based systems are beginning to influence how healthcare is provided.
One significant example is MDAgents, a multi-agent system using LLMs to manage complex medical decisions [25]. Its design replicates the teamwork observed in actual healthcare environments, enabling effective communication among its agents. Testing has shown that MDAgents performs better than earlier models in various evaluations.
A recent review explored the use of LLM-based agents in medical contexts [26]. The review covered technical foundations, practical applications, and existing challenges. It emphasized components such as planning techniques, reasoning strategies, integration of external tools, and agent architecture. These systems are now employed for tasks like clinical decision support, automatic patient documentation, simulation training, and workflow optimization.
Researchers are now further developing LLM-based agents into multi-agent systems (MAS), where multiple agents interact in a decentralized and collaborative way. This change allows for systems that are more organized and flexible, offering new ways to manage challenging healthcare situations, such as emergency response coordination and personalized patient treatment.
@@ II-D Multi-Agent Systems in Healthcare
II-E Ethical Governance in Healthcare AI
@@ II-F Motivation and Research Gap
While clinical decision support systems, LLM-powered agents and multi-agent frameworks [36], have made notable strides in other fields, there’s still a considerable gap when it comes to integrating these technologies into healthcare, especially for real-world ICU settings under clear ethical oversight. Many current solutions fall short in modularity, lack transparency, or aren’t built with the kind of structured, inter-agent communication needed to reflect the fast-paced, interdisciplinary nature of intensive care.
@@ III-A Dataset and Preprocessing
This study used the eICU Collaborative Research Database v2.0 [21], which compiles anonymized ICU records from over 200,000 patient admissions in diverse U.S. hospitals. The database encompasses structured details (e.g., vital signs and lab results) and unstructured clinical notes contributed by nurses and physicians, giving us a thorough view of patient care.
We concentrated on several key eICU files: patient.csv, lab.csv, vitalPeriodic.csv, note.csv, and medication.csv, along with APACHE-related data files [37] (apacheApsVar.csv and apachePatientResult.csv). To align each patient’s information, records were grouped according to patient-unit-stayid. If any essential data were missing—such as vital signs, lab values, or clinical notes—we removed those entries to maintain reliability.
We addressed the data gaps, ordered events based on their timestamps, and shortened lengthy text fields to meet language model guidelines. Then, we sampled 150 patients for the study: 76 mortality patients and 74 survived patients.
Next, we extracted specific features from each patient’s record set. We captured the ten most recent vital sign readings to reveal each patient’s current physiological state. We also selected the latest distinct lab biomarkers deemed clinically relevant. When dealing with unstructured clinical documentation, we included up to three notes for every patient, focusing largely on entries written by physicians and nurses.
Our analysis also tracked the top 20 medications and treatments, identifying them by frequency or uniqueness within the overall dataset. Finally, we incorporated APACHE scores and predictions as reference points to aid in validating and evaluating our modeling outcomes.
@@ III-B Multi-Agent Framework Design
  • Vitals Analysis Agent: Processes vital signs (e.g., heart rate (HR), systolic blood pressure (SBP), peripheral capillary oxygen saturation ($SpO_{2}$), temperature) and evaluates physiological stability, respiratory function, and cardiovascular performance.
@@
To maximize information flow between agents, we implemented a shared memory architecture that allows any agent to access inputs and outputs from previous pipeline stages. This approach mitigates the risk of information loss between modules while maintaining the semantic separation of responsibilities. While this shared memory design has proven effective at improving predictive performance by ensuring that no critical clinical insights are lost during the analysis process, we recognize that the multi-agent architecture introduces additional complexity that can impact transparency. Our current implementation focuses on performance optimization, with ongoing work to enhance the explainability mechanisms across the agent communication pathways.
Each agent is implemented using OpenAI GPT-4o [38] and configured via the intelli framework [39], which allows asynchronous agent orchestration with JSON-structured prompts and logging.
@@ III-D System Execution and DAG Orchestration
Execution is managed asynchronously using Python’s asyncio to allow concurrent LLM calls and reduce latency. The system supports multi-threaded batch evaluation and error-tolerant retries.
Figure 2: Comparison of performance between the Multi-agent and Single-agent frameworks across three evaluation metrics: Mortality Prediction Accuracy, Length of Stay (LOS) Mean Error, and LOS Median Error. Each model was executed 8 times, and the box plots represent the distribution of results over these runs. The Multi-agent framework shows slightly higher mean mortality accuracy with slightly more variance, while LOS-related errors are nearly identical between the two models. These results indicate comparable predictive performance, with minor differences in consistency and central tendency.
III-E Implementation Details
Each agent is instantiated as a generative pre-trained transformer-based (GPT-based) [40] text agent with a predefined mission, API credentials, and output format enforcement. Prompts are customized per task using structured sections (e.g., “KEY ABNORMALITIES”, “APACHE RELEVANT FINDINGS”). Inputs are truncated or summarized to fit within the 10,000-token limit of GPT.
@@ III-E Implementation Details
  model_params={"key": OPENAI_API_KEY,
  "model": "gpt-4o"}
@@ III-F Ethical AI and Explainability
III-G Quantitative Evaluation of Agent Transparency
Here we built the transparency assessment module [41] to evaluate the transparency of clinical prediction explanations by analyzing text responses for key transparency features. It calculates a transparency score by checking for the presence of five critical elements: explicit weights, monotonic relationships, feature importance, decision path explanations, and uncertainty quantification. The scoring process uses regular expression pattern matching to detect relevant terms in the response content. The final transparency score is normalized by dividing the raw score (number of detected features) by the maximum possible score (total number of transparency features), resulting in a value between 0 and 1 that quantifies how transparent and interpretable a clinical prediction explanation is.
@@ IV Results
The experimental results in Table I and Figure 2 demonstrate that the multi-agent system consistently outperforms the single-agent approach across all evaluated metrics over 150 unique patients. Each experiment was conducted across eight runs with approximately 150 patients per run, and the reported values represent the average performance to ensure consistency and robustness of evaluation. In terms of mortality prediction accuracy, the multi-agent model achieved a mean of 59%, while the single-agent model reached only 56%. This 3 percentage point improvement is consistent across multiple runs and represents enhancement in predictive capability in the high-stakes ICU environment. Additionally, the standard deviation for the multi-agent model is marginally higher, suggesting a bit more variability across runs.
TABLE I: Comparison of Multi-agent and Single-agent Models (Average over 8 runs)
@@
Mortality Prediction Accuracy (%) Multi-agent 59.00
Single-agent 56.00
LOS Mean Error (days) Multi-agent 4.37
Single-agent 5.82
Mean Squared Error (days²) Multi-agent 35.49
Single-agent 48.13
Root Mean Squared Error (days) Multi-agent 5.95
Single-agent 6.94
Our analysis indicates an enhancement in predicting Length of Stay (LOS) when using the multi-agent approach. In our study, the average prediction error drops to 4.37 days under the multi-agent strategy, compared to 5.82 days observed with the single-agent method—an improvement of roughly 25% in accuracy. This gain is important, given its direct influence on how ICU resources are allocated and care is planned. In addition, the multi-agent system registers a mean squared error of 35.49 as opposed to 48.13, along with a root mean squared error of 5.95 compared to 6.94, signifying that it delivers more stable and consistent predictions with fewer extreme fluctuations.
A particularly aspect of our findings is how the multi-agent approach manages to reduce the LOS prediction error. A mean error of 4.37 days, as opposed to 5.82 days from the single-agent method, illustrates a noteworthy improvement of 25%, which is critical in fine-tuning patient care. Moreover, the improved metrics—lower mean squared error (35.49 instead of 48.13) and root mean squared error (5.95 compared to 6.94)—further confirm that this system not only enhances accuracy but also offers more stable predictions across various patient groups.
Figure 3: Comparison of Multi-agent and Single-agent frameworks across two key metrics: Mortality Prediction Accuracy (top) and Length of Stay Mean Error (bottom).
Figure 3 reveals a clear analysis across the eight test runs we conducted. Looking at the top graph, you can see the Multi-agent system (blue line) consistently outperformed the Single-agent approach (green line) in predicting mortality. The Multi-agent accuracy ranges from about 56% to nearly 60%, while the Single-agent stays between 55%–57%. What’s notable not just that it performed better, but that this advantage held steady across every single run.
The LOS prediction results in the bottom graph show more improvement for multi-agent model. The blue line stays below the green throughout all runs, with errors around 4–4.7 days compared to the Single-agent’s 5.7–6 days. That gap - somewhere around a day and a half - might not sound huge until you consider what it means for real patients and hospital planning. Most systems like this show more ups and downs, but here the multi-agent setup maintained its edge consistently.
@@
TABLE II: Comparison of transparency score of Multi-agent and Single-agent Models (Average over 8 runs)
Metric Model Mean
Average transparency score (%) Multi-agent 85.50
Single-agent 86.21
Table II provides a side-by-side comparison of average transparency scores from eight independent runs. The data shows that the single-agent approach scores an average of 86.21% in contrast to 85.50% for the multi-agent model. This suggests that, under our current evaluation criteria, both models perform similarly in terms of transparency with the single-agent design being marginally more interpretable. Despite the distributed nature of the multi-agent system, it maintains nearly equivalent transparency levels, indicating that our shared memory architecture effectively preserves reasoning traceability across multiple specialized agents.
@@
Overall, both the multi-agent and single-agent models perform similarly in terms of predictive accuracy and error, with only slight differences that do not strongly favor one over the other. Interestingly, the single-agent model shows a marginal advantage, which might be linked to its avoidance of the complexities that come with managing interactions among multiple agents. Without effective coordination, a multi-agent system may not fully benefit from its distributed learning structure, making centralized processing by a single agent a more effective option under these circumstances. This finding highlights that, in the absence of well-optimized inter-agent dynamics, the simpler single-agent architecture can sometimes outperform a more complex system. Therefore, unless additional factors—such as scalability or deployment constraints—demand a multi-agent setup, both models are viable choices based on the results observed.
@@ V Conclusion
In our recent analysis, we found that the multi-agent system delivers more accurate predictions than the single-agent model. For example, it not only improves mortality prediction rates and shortens the forecasted length of stay—but it does so with comparable interpretability. Specifically, the single-agent model scores 86.21% for transparency, whereas the multi-agent system comes in at a very close 85.50%. This minimal difference suggests that our multi-agent architecture effectively maintains transparency despite the inherent complexity of coordinating decisions among several specialized agents.
This balance between performance and understandability has serious implications in the clinical arena. In scenarios where prediction accuracy is essential—such as in critical care decisions—the multi-agent approach clearly offers benefits. Conversely, when it is crucial to explain the reasoning behind predictions to build clinician trust, the higher transparency of the single-agent model is more attractive.
Looking ahead, our research will concentrate on boosting the clarity of the multi-agent framework without compromising its predictive strengths. By enhancing inter-agent communication protocols and incorporating more advanced explanation mechanisms, we aim to develop a system that marries the superior predictive power of multi-agent architectures with the interpretability required for safe and effective use in critical care.